OpenStudy (anonymous):

How do I integrate x(e^-x/2) dx

OpenStudy (tkhunny):

?? Is it \(x^{2}\;or\;\dfrac{x}{2}\)?

OpenStudy (anonymous):

e to the -x/2 power

OpenStudy (primeralph):

use integration by parts

zepdrix (zepdrix):

\[\large \int\limits xe^{-x/2}\;dx\]

OpenStudy (anonymous):

when I use by parts it comes down to \[\frac{e ^{-x/2} x^2 }{ 2} + \frac{ 1 }{ 4 } \int\limits x^2 e ^{-x/2} dx\]

OpenStudy (primeralph):

switch your u and v

OpenStudy (anonymous):

got it thank you (: